The common stock of Auto Deliveries sells for $28.16 a share. The stock is expected to...

  1. The common stock of Auto Deliveries sells for $28.16 a share. The stock is expected to pay $1.35 per share next year when the annual dividend is distributed. The firm has established a pattern of increasing its dividends by 3% annually and expects to continue doing so. What is the required rate of return on this stock?

    A.

    7.42%

    B.

    7.79%

    C.

    19.67%

    D.

    20.14%

    E.

    20.86%
Answer #1

Req Ret = [ Expected Div / Price ] + Growth rate

= [ 1.35 / 28.16 ] + 0.03

= 0.0479 + 0.03

= 0.0779

i.e 7.79%

Option B is correct.
Answer #2

Given :


P0 = 28.16 ($)

D1 = 1.35 ($)

g = 3 % = 0.03



Now,


Market rate of return, r 

= D1 / P 0 + g

= 1.35 / 28.16 + 0.03

= 0.0779

= 7.79 %



Market rate of return is : Option B.   7.79%  (ANSWER). 


answered by: Tulsiram Garg